OpenStudy (anonymous):

Extremal points, local/global maxima/minima f(x) = x^2 + px + q, x E (-infinity, infinity)

OpenStudy (anonymous):

I first made f'(x) = 0 so.. 2x+ p = 0, so x = -1/2 p but I have no idea what to do next :S help me please

OpenStudy (anonymous):

f''(x) = 2 so its a minimum, but how do I tell if its global or local?????

OpenStudy (anonymous):

minima which is smallest among all minimas then its global

OpenStudy (anonymous):

but how did you know it was global?

OpenStudy (anonymous):

have u find all the minimas of the equation

OpenStudy (anonymous):

All I have really done is do the 2nd Derivation to find if it was maxima or minima but I don't know how to work out if it is local or global

OpenStudy (anonymous):

if there is only one minima then its global ,

OpenStudy (anonymous):

local minima is a relative term

OpenStudy (anonymous):

I don't know if there is only one minima, thats what I really need help with :S

OpenStudy (anonymous):

:)its tricky

OpenStudy (anonymous):

so for f(x) = x^2 + px + q, I make f'(x) = 0 so x = -1/2p, what do I do after that??

OpenStudy (anonymous):

dont know for diffrent values of x we will have diffrent minima ,as x depends on p ,

OpenStudy (loser66):

f(x) = x^2 +px +q is an upward parabola ---> it has minimum point only, right?

OpenStudy (anonymous):

yes u r right bt the parabola cn be anywhere depends on the value of discremenent

OpenStudy (loser66):

and its vertex (-p/2, q) ok?

OpenStudy (loser66):

if p, q are arbitrary, so?? just argue on them. We don't need a special points to argue on, right?

OpenStudy (loser66):

The question is about local/ global max, min Our parabola is upward, then we have min only,

OpenStudy (loser66):

the form of the parabola shows that it doesn't have local, then that min is global

OpenStudy (anonymous):

we have min bt there cn be a no of minimas

OpenStudy (loser66):

and the parabola gets global min at -p/2 and its value is q And we are done. That's all I know. :)

OpenStudy (anonymous):

sorry for being slow, but can you explain why its value is q? :S

OpenStudy (loser66):

the form \(y = ax^2+bx+c\) has vertex at \(x =\dfrac{-b}{2a}\) to this problem \(x = \dfrac{-p}{2}\) , then plug it back to find y of the vertex \((\dfrac{-p}{2})^2 +p*(\dfrac{-p}{2})+q\) \(\dfrac{p^2}{4}-\dfrac{p^2}{4}+q=q\) so that vertex is (p/2,q)

OpenStudy (loser66):

oh, I am sorry, mistake :)

OpenStudy (anonymous):

isnt it p^2/4 - p^2/2 ?

OpenStudy (loser66):

yes, I am sooooo sorry for my carelessness. :)

OpenStudy (anonymous):

no problem I appreciate you helping :)

OpenStudy (loser66):

But ganeshie8 is here, he can handle it. right? @ganeshie8 rescue me, pleeeeeease

OpenStudy (anonymous):

would it be, global minimum at point (-1/2p , q - 3/4p^2)?

ganeshie8 (ganeshie8):

try this instead http://www.wolframalpha.com/input/?i=min+and+max+x%5E2+%2B+c*x+%2B+d

OpenStudy (anonymous):

could you please explain how to tell if its global or local please?

ganeshie8 (ganeshie8):

it is a quadratic, so you can have atmost one extrema

ganeshie8 (ganeshie8):

since the leading coefficient is positive its a upward facing parabola 

ganeshie8 (ganeshie8):

so thats the only extrema and it is global too

OpenStudy (anonymous):

makes sense, thank you :) , another example is min max, local or global for xlnx where x E [1,10] is that the sample idea as the example above?

ganeshie8 (ganeshie8):

yes, but u need to be careful when you're given an interval

ganeshie8 (ganeshie8):

the extrema can occur at boundary also

ganeshie8 (ganeshie8):

you need to evaluate the function at x = 1 and x = 10 also and compare them against other values at critical points

ganeshie8 (ganeshie8):

thats for global min/max ^^

OpenStudy (anonymous):

so xlnx will have a min and max? because its not quadratic? will there be local and global ones also?

ganeshie8 (ganeshie8):

we don't know that yet, lets find the critical points by solving f' = 0

ganeshie8 (ganeshie8):

find the derivative, set it equal to 0 and solve x

OpenStudy (anonymous):

f'(x) = 1 + ln(x) so 1+ln(x) = 0, so x = 1/e ?

ganeshie8 (ganeshie8):

yup! so ur critical x values are : x = 1/e x = 1 x = 10

OpenStudy (anonymous):

so f(1/e) = 1/e * ln(1/e) = -1/e?

ganeshie8 (ganeshie8):

yes! evaluate the function at boundary values too

ganeshie8 (ganeshie8):

x = 1/e, y = -1/e x = 1, y = ? x = 10, y = ?

OpenStudy (anonymous):

oh I see one sec